Reynoso Corporation manufactures titanium and aluminum tennis racquets. Reynoso’s total overhead costs consist of assembly costs and inspection costs. The following information is available:

Cost                      Titaniium      Aluminum     Total Cost

Assembly                 500 mach. hours    500 mach. hours $60,000
Inspections      350 250 $100,000
                                   2,100 labor hours    1,900 labor hours


Reynoso is considering switching from one overhead rate based on labor hours to activity-based costing.
Total overhead costs assigned to titanium racquets, using a single overhead rate, are

$80,000.

$84,000.

$100,000

$112,000.
Answer #1
Concepts and reason

Cost: A monetary term which refers as the valuation of any additional service such as payment for the delivery of the goods and services, and the amount of acquisition of goods is known as cost.

Activity-Based Costing: Activity-based costing is the method to analyze the each activity according to the specified costs. The cost of the activity is defined on the basis of the products and services are consumed by them.

Total cost: The combination of fixed and variable cost is known as the total cost. Companies differentiate between the fixed and variable cost and evaluate both the costs in differently.

Fundamentals

Manufacturing Overheads: The cost which is indirectly related to the production is known as manufacturing overheads.

Fixed overheads: Overheads which remain constant and don’t change due to change in level of production over a period are known as fixed overheads.

Variable overheads: The overheads which don’t remain constant and change as per the production level of the company change. The variable overheads per unit remains constant and the total variable cost changes as per the level of production change.

The formula to calculate the total overheads is shown below.

Totaloverheads=Totalcostofassembly+Totalcostofinspection{\rm{Total}}\;{\rm{overheads}} = {\rm{Total}}\;{\rm{cost}}\;{\rm{of}}\;{\rm{assembly}} + {\rm{Total}}\;{\rm{cost}}\;{\rm{of}}\;{\rm{inspection}}

The formula to calculate the total labor of hours is shown below.

Totallaborsofhours=Laborofhoursofassembly+Laborofhourseofinspection{\rm{Total}}\;{\rm{labors}}\;{\rm{of}}\;{\rm{hours}} = {\rm{Labor}}\;{\rm{of}}\;{\rm{hours}}\;{\rm{of}}\;{\rm{assembly}} + {\rm{Labor}}\;{\rm{of}}\;{\rm{hourse}}\;{\rm{of}}\;{\rm{inspection}}

The formula to calculate the total overheads assigned to titanium is shown below.

Totaloverheadstotitanium=TotaloverheadsTotallaborofhours×Laborhoursoftitanium{\rm{Total}}\;{\rm{overheads}}\;{\rm{to}}\;{\rm{titanium}} = \frac{{{\rm{Total}}\;{\rm{overheads}}}}{{{\rm{Total}}\;{\rm{labor}}\;{\rm{of}}\;{\rm{hours}}}} \times {\rm{Labor}}\;{\rm{hours}}\;{\rm{of}}\;{\rm{titanium}}

Compute the total overheads.

Totaloverheads=Totalcostofassembly+Totalcostofinspection=$60,000+$100,000=$160,000\begin{array}{c}\\{\rm{Total}}\;{\rm{overheads}} = {\rm{Total}}\;{\rm{cost}}\;{\rm{of}}\;{\rm{assembly}} + {\rm{Total}}\;{\rm{cost}}\;{\rm{of}}\;{\rm{inspection}}\\\\ = {\rm{\$ 60,000}} + \$ 100,000\\\\ = \$ 160,000\\\end{array}

Therefore, the total overheads are $160,000.

Compute the total labor of hours.

Totallaborsofhours=(Laborofhoursofassembly+Laborofhourseofinspection)=2,100hours+1,900hours=4,000hours\begin{array}{c}\\{\rm{Total}}\;{\rm{labors}}\;{\rm{of}}\;{\rm{hours}} = \left( \begin{array}{l}\\{\rm{Labor}}\;{\rm{of}}\;{\rm{hours}}\;{\rm{of}}\;{\rm{assembly}}\\\\ + {\rm{Labor}}\;{\rm{of}}\;{\rm{hourse}}\;{\rm{of}}\;{\rm{inspection}}\\\end{array} \right)\\\\ = 2,100\;{\rm{hours}} + 1,900\;{\rm{hours}}\\\\ = 4,000\;{\rm{hours}}\\\end{array}

Therefore, the total labors of hours are 4,000 hours.

Compute the total overheads assigned to titanium racquets.

Totaloverheadstotitanium=TotaloverheadsTotallaborofhours×Laborhoursoftitanium=$160,0004,000hours×2,100hours=$40×2,100=$84,000\begin{array}{c}\\{\rm{Total}}\;{\rm{overheads}}\;{\rm{to}}\;{\rm{titanium}} = \frac{{{\rm{Total}}\;{\rm{overheads}}}}{{{\rm{Total}}\;{\rm{labor}}\;{\rm{of}}\;{\rm{hours}}}} \times {\rm{Labor}}\;{\rm{hours}}\;{\rm{of}}\;{\rm{titanium}}\\\\ = \frac{{\$ 160,000}}{{4,000\;{\rm{hours}}}} \times 2,100\;{\rm{hours}}\\\\ = \$ 40 \times 2,100\\\\ = \$ 84,000\\\end{array}

Therefore, the total overheads assigned to titanium are $84,000.

Ans:

The total overheads assigned to titanium are $84,000.